# A Swift and Omnidirectional Formation Approach based on Hierarchical Reorganization

### v1

@@ III-B Affine Formation Maneuver Control
To simplify the expression of Eq.(1), the stress matrix is defined as $\mathbf{\Omega}\in\mathbf{R}^{n\times n}$ which satisfies
@@
(Stress Condition for Affine Localizability [8]) The nominal formation ($\mathcal{G},\boldsymbol{r}$) is affinely localizable if and only if $\boldsymbol{\bar{\Omega}}_{ff}$ is nonsingular. When $\boldsymbol{\bar{\Omega}}_{ff}$ is nonsingular, for any $\boldsymbol{p}=[\boldsymbol{p}_{l}^{T},\boldsymbol{p}_{f}^{T}]^{T}\in\mathcal{A}(\boldsymbol{r})$, $\boldsymbol{p}_{f}$ can be uniquely calculated as $\boldsymbol{p}_{f}=-\boldsymbol{\bar{\Omega}}_{ff}^{-1}\boldsymbol{\bar{\Omega}}_{fl}\boldsymbol{p}_{l}$.
Figure 1: The architecture of the proposed omnidirectional affine formation strategy. In formations eligible for hierarchical reorganizations, the initial step involves validating individual roles, as determined by a minimum time-driven leadership evaluation algorithm. As the formation nears the transition juncture, followers designated to assume leadership are seamlessly transitioned through implementing a power transition control algorithm. Upon arrival at the transition point, a strategic reorganization in the formation’s topological structure is executed, utilizing a power-centric topology switching strategy. This reconfigured topological structure then serves as the basis for applying the power transition control strategy. This application is focused on the smooth devolution of authority from the leader transitioning to a follower role, thereby ensuring a fluid and uninterrupted transition within the hierarchical reorganization process of the formation.
III-C Problem Statement
During the traversal of an aerial formation through any environment $\mathbf{E}$, the complexity of $\mathbf{E}$ imposes constraints on the traversal direction. Building on this, agile drone formations must be capable of high-maneuverability turns. For omnidirectional affine formations, the traversal direction is determined by the formation hierarchy configuration $\mathcal{G}^{i}(\mathcal{V}^{i},\mathcal{E}^{i}),i=1,2,...,r$. To facilitate the formation’s swift adaptation to any given environment $\mathbf{E}$, this problem can be modeled as follows:
$\mathcal{G}^{i}(\mathcal{V}^{i},\mathcal{E}^{i})=\digamma(^{j}\boldsymbol{x},^{j}\hat{\boldsymbol{x}},\mathbb{G})$ (4)
where $\mathbb{G}=\{\mathcal{G}^{1}(\mathcal{V}^{1},\mathcal{E}^{1}),...,\mathcal{G}^{r}(\mathcal{V}^{r},\mathcal{E}^{r})\}$ is the feasible formation configuration set, the ${}^{j}\boldsymbol{x}$ and ${}^{j}\hat{\boldsymbol{x}}$ represent the current state and target state of drone $j$, respectively.
IV Omnidirectional affine formation based on hierarchical reorganizations
@@
This section comprehensively elaborates on the specific methods of omnidirectional affine formation, divided into three distinct parts, with the overall system framework depicted in Fig. 1. The figure illustrates the inherent connections between these different components. Firstly, we define the omnidirectional affine formation and analyze the necessary conditions for hierarchical reorganizations. On this basis, we establish a minimum time-driven leadership evaluation algorithm tailored for the flexibility and coordination of the swarm to facilitate role recognition among different robots. Given the challenges of synchronous role switching in actual systems, we first introduce a power transition control strategy for smooth power transition by designing a power handover function. To eliminate disturbances caused by other followers, a power-centric topology switching strategy is proposed to facilitate the smooth execution of hierarchical reorganizations.
IV-A Omnidirectional Affine Formation
Figure 2: Illustration of Initial Configuration for Omnidirectional Affine Formation. (a) Four drones formation. (b) Five drones formation.
@@ IV-A Omnidirectional Affine Formation
(Dynamic Affine Localizability) For the changing graph configuration $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$, the position of followers $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{f}(t)}$ can be uniquely determined by the position of leaders $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{l}(t)}$.
@@
(Initial Formation for Dynamic Affine Localizability): For the initial formation $(\mathcal{G}^{o},\boldsymbol{r}^{o})$, assume that $\{\boldsymbol{r}^{o}_{i}\}_{i\in\mathcal{V}^{o}_{l}}$ affinely span $\mathbf{R}^{d}$.
@@
(Hierarchical Reorganization for Dynamic Affine Localizability): Under Assumption 1, the changing formation $(\mathcal{G}^{t},\boldsymbol{r}^{t})$ is dynamically affine localizable if and only if $\{\boldsymbol{r}^{t}_{i}\}_{i\in\mathcal{V}^{t}_{l}}$ affinely span $\mathbf{R}^{d}$.
@@
(Sufficiency) According to the Theorem 1 in [8], the formation $(\mathcal{G},\boldsymbol{r})$ is affinely localizable if and only if $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{l}}$ affinely span $\mathbf{R}^{d}$. For any formation $(\mathcal{G}^{t_{1}},\boldsymbol{r}^{t_{1}})\in(\mathcal{G}^{t},\boldsymbol{r}^{t})$, $(\mathcal{G}^{t_{1}},\boldsymbol{r}^{t_{1}})$ is affinely localizable if $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{l}}^{t_{1}}$ affinely span $\mathbf{R}^{d}$. Then, $(\mathcal{G}^{t},\boldsymbol{r}^{t})$ is dynamically affine localizable consequently.
(Necessity) For any formation $(\mathcal{G}^{t_{1}},\boldsymbol{r}^{t_{1}})\in(\mathcal{G}^{t},\boldsymbol{r}^{t})$, $(\mathcal{G}^{t_{1}},\boldsymbol{r}^{t_{1}})$ cannot realize affine localizability if $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{l}}^{t_{1}}$ cannot affinely span $\mathbf{R}^{d}$. Then, $(\mathcal{G}^{t},\boldsymbol{r}^{t})$ cannot satisfy dynamically affine localizability consequently. ∎
@@
(Omnidirectional Affine Formation) For the changing formation $(\mathcal{G}^{t},\boldsymbol{r}^{t})$, any target configuration $\boldsymbol{p}^{*}$ of current nominal configuration $\boldsymbol{r}^{t}$ locates at the same affine image $\mathcal{A}(\boldsymbol{r}^{t})$.
@@
(Initial Configuration for Omnidirectional Affine Formation) For any initial configuration $\boldsymbol{r}^{o}$, assume that $C(n,n_{l})\geq 1$ where $C(n,n_{l})$ represents the count of viable combinations for hierarchical reorganization within $\boldsymbol{r}^{o}$.
@@
(Omnidirectional Affine Formable) Under Assumptions 1 and 2, the dynamic configuration $\boldsymbol{p}$ is omnidirectional affine formable if and only if the initial configuration $\boldsymbol{r}^{o}$ satisfies that $\{\boldsymbol{r}^{o_{j}}\}_{j\in[1,C(n,n_{l})]}\in\mathcal{A}(\boldsymbol{r}^{o})$.
@@
Based on the property of affine transformation, the matrix $\boldsymbol{A}$ is invertible [33], which means the inverse of $\boldsymbol{I}_{n}\otimes\boldsymbol{A}^{\prime}$ exists. Then,
@@
Several omnidirectional affine formations are depicted in Fig.2. For a changing formation $\mathcal{G}^{t}$ satisfying Theorem IV.1 and IV.2, we can identify a collection of viable formations from $\{\boldsymbol{r}^{o_{j}}\}_{j\in[1,C(n,n_{l})]}\in\mathcal{A}(\boldsymbol{r}^{o})$ represented as $\mathbb{G}$.
Figure 3: Process of leadership evaluation.
In determining the optimal formation post-hierarchical reorganization, the leadership evaluation encompasses a twofold assessment: the overall flexibility of the formation and the maneuverability of each agent. Fig. 3 illustrates the minimum time-driven leadership evaluation algorithm process, providing a clear visual representation of the methodology applied to hierarchical reorganization. First, when the relevant formation collection $\mathcal{G}^{i}$ is determined, the leadership evaluation index can be initially defined as:
$\Theta\leftarrow\min_{i=1,2,...,r}(t^{\mathcal{G}^{i}})\quad\mathcal{G}^{i}\in\mathbb{G}$ (14)
where $t^{\mathcal{G}^{i}}$ represents the required time for the formation collection $\mathcal{G}^{i}$ from current location to target location. Specifically, the $t^{\mathcal{G}^{i}}_{j}$ is expressed as the flight time required for the agent $j$ from its current location ${}^{j}\boldsymbol{\hat{x}}$ to reach its corresponding target location in $\mathcal{G}^{i}$, which is denoted as ${}^{j}\boldsymbol{\hat{x}}^{G^{i}}$. In this paper, the $\Phi(\cdot)$ representing the trajectory generation algorithm is applied to calculate $t^{\mathcal{G}^{i}}_{j}$ and obtain the set of trajectory points $\mathbb{P}^{j}$ of agent $j$, which is formulated as:
$(t^{\mathcal{G}^{i}}_{j},\mathbb{P}^{j})=\Phi(^{j}\boldsymbol{\hat{x}},^{j}\boldsymbol{\hat{x}}^{\mathcal{G}^{i}})$ (15)
$s.t.\quad\max(^{j}\boldsymbol{x}^{(1)})\leq^{j}v^{max}\quad\min(^{j}\boldsymbol{x}^{(1)})\geq^{j}v^{min}$
$\max(^{j}\boldsymbol{x}^{(2)})\leq^{j}a^{max}\quad\min(^{j}\boldsymbol{x}^{(2)})\geq^{j}a^{min}$
where ${}^{j}v^{max}$ and ${}^{j}v^{min}$ represent the maximum and minimum velocities for agent $j$, respectively. Meanwhile, ${}^{j}a^{max}$ and ${}^{j}a^{min}$ represent the maximum and minimum accelerations for agent $j$, respectively. To ensure the maneuverability of each agent while achieving synchronicity within the formation. For any agent $j$, the flight time is standardized to the maximum value of the individual planning time. Then, for the formation collection $\mathbb{G}$, the leadership evaluation index can be reformulated as:
$(\Theta,\tau)\leftarrow\min_{i=1,2,...,r}(\max_{k=1,2,...,n}(t_{k}^{\mathcal{G}^{i}}))$ (16)
$s.t.\quad\mathcal{G}^{i}\in\mathbb{G}$
where $\tau$ represents the actual planning time with the selected formation configuration $\Theta$.
IV-B Power Transition Control Algorithm
In hierarchical formation, control under affine transformations typically manifests as leaders and followers employing distinct strategies. We first consider the control of the leader. Based on the formation configuration $\Theta$ and the planned time $\tau$ derived from Eq.(16), the set of discrete point motion states for each agent $i$ at each moment can be expressed as:
$\mathbb{P}^{i}=\Phi(^{i}\boldsymbol{\hat{x}},^{i}\boldsymbol{\hat{x}}^{\Theta},\tau)$ (17)
Therefore, for the leader $i$ at time $T(k+1)$, the target position is $\boldsymbol{\hat{p}}_{i}^{T(k+1)}\in\mathbb{P}^{i}$. For the leader, it is only necessary to track the state of the target point in real-time at $T(k+1)$.
Secondly, the control strategy of follower $i$ at time $T(k+1)$ is described as:
$\boldsymbol{p}^{T(k+1)}_{i}=\boldsymbol{p}^{T(k)}_{i}-\frac{\sum_{j\in N_{i}}\omega_{i}^{j}(\boldsymbol{p}_{i}^{T(k)}-\boldsymbol{p}_{j}^{T(k)}-\boldsymbol{v}_{j}^{T(k)})}{\sum_{j\in N_{i}}\omega_{i}^{j}}\delta T(k+1)$ (18)
$\delta T(k+1)=T(k+1)-T(k)$
This strategy achieves the control of the follower, maintaining the affine formation while following the leader.
Unlike the control strategy, where the leader does not switch, hierarchical reorganization reflects the power transfer. It is evident that Eq.(17) represents the pinnacle of individual power. Conversely, Eq.(18) signifies the complete dissipation of individual power. To address this characteristic, we design a power transfer function to demonstrate the features of different power transfer processes. Since the transition from follower to leader signifies an expansion of individual rights, and the transition from leader to follower represents a reduction of individual power, the power transfer function is designed as follows:
[…]
Where $t\in[0,\tau]$. The power transfer function is shown in Fig. 4.
Figure 4: The power transfer function.
Consequently, during the hierarchical reorganization process, the control function can be designed as:
$\boldsymbol{{}^{d}p}^{T(k+1)}_{i}=(1-\gamma(t))\boldsymbol{p}^{T(k+1)}_{i}+\gamma(t)\boldsymbol{\hat{p}}_{i}^{T(k+1)}$ (20)
Assumption 3.
(Target Trajectory for Followers) For any follower $i$, assume that $\boldsymbol{p}^{T(k+1)}_{i}$ is on the target trajectory.
Theorem IV.3.
(Convergence of Power Transition Control) Under the Assumption 3, the convergence rate of Eq. (20) is faster than $e^{t}$.
Proof.
According to the definition of $\boldsymbol{\hat{p}}_{i}^{T(k+1)}$, the control error can be expressed as:
$\boldsymbol{{}^{d}\delta p}^{T(k+1)}_{i}$ $=\boldsymbol{\hat{p}}_{i}^{T(k+1)}-[(1-\gamma(t))\boldsymbol{p}^{T(k+1)}_{i}+\gamma(t)\boldsymbol{\hat{p}}_{i}^{T(k+1)}]$ (21)
$=(1-\gamma(t))(\boldsymbol{\hat{p}}_{i}^{T(k+1)}-\boldsymbol{p}^{T(k+1)}_{i})$
Denote $\boldsymbol{\delta p}^{T(k+1)}_{i}=\boldsymbol{\hat{p}}_{i}^{T(k+1)}-\boldsymbol{p}^{T(k+1)}_{i}$, then the control error is reformulated as:
$\boldsymbol{{}^{d}\delta p}^{T(k+1)}_{i}=(1-\gamma(t))(\boldsymbol{\delta p}^{T(k+1)}_{i})$ (22)
Substitute $\boldsymbol{p}^{T(k+1)}_{i}$ with Eq.(18), we can obtain that
$\boldsymbol{\delta p}^{T(k+1)}_{i}=\boldsymbol{\hat{p}}_{i}^{T(k+1)}-\boldsymbol{p}^{T(k+1)}_{i}=\boldsymbol{\hat{p}}_{i}^{T(k+1)}-\boldsymbol{p}^{T(k)}_{i}+\frac{1}{\sum_{j\in N_{i}}\omega_{i}^{j}}$ (23)
$(\sum_{j\in N_{i}}\omega_{i}^{j}(\boldsymbol{p}_{i}^{T(k)}-\boldsymbol{p}_{j}^{T(k)}-\boldsymbol{v}_{j}^{T(k)}))\delta T(k+1)$
Simplify the Eq.(23), we can obtain that:
$\sum_{j\in N_{i}}\omega_{i}^{j}(\boldsymbol{p}_{i}^{T(k)}-\boldsymbol{p}_{j}^{T(k)})+\sum_{j\in N_{i}}\omega_{i}^{j}(\boldsymbol{v}_{i}^{T(k)}-\boldsymbol{v}_{j}^{T(k)})=0$ (24)
According to Section III-B, the local matrix-vector form in Eq.(24) is
$\sum_{j\in N_{i}}\omega_{i}^{j}(\boldsymbol{p}_{i}^{T(k)}-\boldsymbol{p}_{j}^{T(k)})$ $=\boldsymbol{\bar{\Omega}_{ff}p_{f}}+\boldsymbol{\bar{\Omega}_{fl}p^{*}_{l}}=\boldsymbol{\bar{\Omega}_{ff}\delta p_{f}}$ (25)
$\sum_{j\in N_{i}}\omega_{i}^{j}(\boldsymbol{v}_{i}^{T(k)}-\boldsymbol{v}_{j}^{T(k)})$ $=\boldsymbol{\bar{\Omega}_{ff}v_{f}}+\boldsymbol{\bar{\Omega}_{fl}v^{*}_{l}}=\boldsymbol{\bar{\Omega}_{ff}\delta v_{f}}$
Then, the matrix-vector of Eq.(24) can be formulated as:
$\boldsymbol{\bar{\Omega}_{ff}\delta p_{f}}+\boldsymbol{\bar{\Omega}_{ff}\delta v_{f}}=\mathbf{0}$ (26)
As $\boldsymbol{\delta v_{f}}=\boldsymbol{\dot{\delta}p_{f}}$ and $\boldsymbol{\bar{\Omega}_{ff}}$ is positive definite, we have $\boldsymbol{\dot{\delta}p}=\boldsymbol{\delta p}$, which implies $\boldsymbol{\delta p}$ converges to zero globally and exponentially fast.
According to the definition of $\gamma(t)$, $\gamma(t)\leq 1$ always holds, which means $1-\gamma(t)\leq 1$ always satisfies. From Eq.(22), the control error satisfies:
$\boldsymbol{{}^{d}\delta p}^{T(k+1)}_{i}=(1-\gamma(t))\boldsymbol{\delta p}^{T(k+1)}_{i}\leq\boldsymbol{\delta p}^{T(k+1)}_{i}$ (27)
which implies that the proposed control function converges globally and at a faster rate than $e^{t}$. ∎
IV-C Power-Centric Topology Switching
Figure 5: Illustration for power-centric topology switching.

### v2

@@ III-C Stress Matrices and Affine Maneuver
To simplify the expression of Eq.(1), the stress matrix is defined as $\mathbf{\Omega}\in\mathbb{R}^{n\times n}$ which satisfies
@@
The nominal formation ($\mathcal{G},\boldsymbol{r}$) is affinely localizable if and only if $\boldsymbol{\bar{\Omega}}_{ff}$ is nonsingular. When $\boldsymbol{\bar{\Omega}}_{ff}$ is nonsingular, for any $\boldsymbol{p}=[\boldsymbol{p}_{l}^{T},\boldsymbol{p}_{f}^{T}]^{T}\in\mathcal{A}(\boldsymbol{r})$, $\boldsymbol{p}_{f}$ can be uniquely calculated as $\boldsymbol{p}_{f}=-\boldsymbol{\bar{\Omega}}_{ff}^{-1}\boldsymbol{\bar{\Omega}}_{fl}\boldsymbol{p}_{l}$.
III-D Problem Statement
During the process of a formation navigating through any given environment $E$, the formation shape is often constrained by the characteristics of the environment $E$. For a reconfigurable hierarchical formation $(\mathcal{G}^{t},\boldsymbol{r}^{t})$, dynamic reorganization of the leader layer typically implies different travel advantages under the environment $E$. These advantages may include the shortest travel time, optimal formation visibility, the most stable communication, and so on. Building on this, we construct an optimization equation according to the travel advantages associated with formation reorganization:
Figure 1: A.The system architecture of reconfigurable hierarchical formation. B.Comparison of with leader-follower affine formation(LFA)[Zhao2018], virtual structure formation(VSF)[Zhou2018], dynamic leader selection(DLS)[Li2017], particle swarm optimization(PSO)[Zhou2022] and our proposed method RHF. The radar chart shows that moving outward from the center represents increasing corresponding versatile. Specifically, stability represents the ability of the formation to maintain its standard shape during movement. Resilience represents the formation’s ability to recover when individual agents lose connectivity. Flexibility represents the formation’s adaptability to complex environments. Optimality, as defined in [Zhou2022], represents the ability to seek optimal formation in spatial, temporal, or other user-defined scales. Moreover, extensibility represents the ability to analyze and model formations for specific tasks.
$J=\phi(\mathcal{G}^{t},\boldsymbol{r}^{t},E)$ (4)
$s.t.\quad det(\boldsymbol{\Omega}_{ff}^{\mathcal{G}^{t}})>0\quad\quad\boldsymbol{r}^{t}\in\mathcal{A}(\boldsymbol{r}^{o})$
where $\phi(\cdot)$ denotes the function that represents different maneuvering advantages under the current environmental conditions $E$. The condition $det(\Omega_{ff})>0$ represents the topology that must be satisfied to allow the positions of the follower layer to be determined by their neighbors. Meanwhile, $\boldsymbol{r}^{t}\in\mathcal{A}(\boldsymbol{r}^{o})$ specifies the positional constraints to guarantee the feasibility of the formation reorganization. This paper focuses on analyzing the constraints of the proposed optimization equation, aiming to develop a reconfigurable hierarchical formation that is both flexible and resilient.
IV Feasible Conditions For Reconfigurable Hierarchical Formation
@@
This section comprehensively elaborates on the specific methods of reconfigurable hierarchical formation, divided into two distinct parts, including the two parts of hierarchical reorganization. Firstly, we define the reconfigurable hierarchical formation and analyze the necessary conditions for hierarchical reorganizations. To eliminate disturbances caused by other followers, a power-centric topology switching strategy is proposed to facilitate the smooth execution of hierarchical reorganizations.
IV-A Reconfigurable Hierarchical Formation
Figure 2: Illustration of Feasible Configurations for Reconfigurable Hierarchical Formation.
Dynamic affine localizability, designed to accommodate hierarchical reconfiguration of the formation, means that the roles of agents within the formation can switch between leader and follower. Let the changing graph $\mathcal{G}^{t}=\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$, we define dynamic affine localization as:
@@ IV-A Reconfigurable Hierarchical Formation
(Dynamic Affine Localizability) For the changing formation $(\mathcal{G}^{t},\boldsymbol{r}^{t})$, the position of followers $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{f}(t)}$ can be uniquely determined by the position of leaders $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{l}(t)}$.
@@
(Initial Formation for Dynamic Affine Localizability): For the initial formation $(\mathcal{G}^{o},\boldsymbol{r}^{o})$, assume that $\{\boldsymbol{r}^{o}_{i}\}_{i\in\mathcal{V}^{o}_{l}}$ affinely span $\mathbb{R}^{d}$.
@@
(Hierarchical Reorganization for Dynamic Affine Localizability): Under Assumption 1, the changing formation $(\mathcal{G}^{t},\boldsymbol{r}^{t})$ is dynamically affine localizable if and only if $\{\boldsymbol{r}^{t}_{i}\}_{i\in\mathcal{V}^{t}_{l}}$ affinely span $\mathbb{R}^{d}$.
@@
(Sufficiency) According to Lemma III.1, the formation $(\mathcal{G},\boldsymbol{r})$ is affinely localizable if and only if $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{l}}$ affinely span $\mathbb{R}^{d}$. For any formation $(\mathcal{G}^{t_{1}},\boldsymbol{r}^{t_{1}})\in(\mathcal{G}^{t},\boldsymbol{r}^{t})$, $(\mathcal{G}^{t_{1}},\boldsymbol{r}^{t_{1}})$ is affinely localizable if $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{l}}^{t_{1}}$ affinely span $\mathbb{R}^{d}$. Then, $(\mathcal{G}^{t},\boldsymbol{r}^{t})$ is dynamically affine localizable consequently.
(Necessity) For any formation $(\mathcal{G}^{t_{1}},\boldsymbol{r}^{t_{1}})\in(\mathcal{G}^{t},\boldsymbol{r}^{t})$, $(\mathcal{G}^{t_{1}},\boldsymbol{r}^{t_{1}})$ cannot realize affine localizability if $\{\boldsymbol{r}_{i}\}_{i\in\mathcal{V}_{l}}^{t_{1}}$ cannot affinely span $\mathbb{R}^{d}$. Then, $(\mathcal{G}^{t},\boldsymbol{r}^{t})$ cannot satisfy dynamically affine localizability consequently. ∎
Formations can better leverage the advantages of the group through hierarchical reorganization. We call formations with this characteristic a reconfigurable hierarchical formation, which is defined as:
@@
(Reconfigurable hierarchical formation) For the changing formation $(\mathcal{G}^{t},\boldsymbol{r}^{t})$, any target configuration $\boldsymbol{p}^{*}$ of current nominal configuration $\boldsymbol{r}^{t}$ locates at the same affine image $\mathcal{A}(\boldsymbol{r}^{t})$.
@@
(Initial Configuration for Versatile Affine Formation) For any initial configuration $\boldsymbol{r}^{o}$, assume that $C(n,n_{l})\geq 1$ where $C(n,n_{l})$ counts the number of viable combinations for hierarchical reorganization.
@@
(Reconfigurable Affine Formable) Under Assumptions 1 and 2, the dynamic configuration $\boldsymbol{r}^{t}$ is reconfigurable affine formable if and only if the initial configuration $\boldsymbol{r}^{o}$ satisfies that $\{\boldsymbol{r}^{o_{j}}\}_{j\in[1,C(n,n_{l})]}\in\mathcal{A}(\boldsymbol{r}^{o})$.
@@
Based on the property of affine transformation, the matrix $\boldsymbol{A}$ is invertible [Lin2016], which means the inverse of $\boldsymbol{I}_{n}\otimes\boldsymbol{A}^{\prime}$ exists. Then,
@@
For a changing formation $\mathcal{G}^{t}$ satisfying Theorem IV.1 and IV.2, we can identify a collection of viable formations from $\{\boldsymbol{r}^{o_{j}}\}_{j\in[1,C(n,n_{l})]}\in\mathcal{A}(\boldsymbol{r}^{o})$ represented as $\mathbf{r}$.
IV-B Power-Centric Topology Switching
Figure 3: Examples for 2d and 3d dynamic graph configurations.